OpenStudy (anonymous):

Suppose every edge of a cube measures centimeters. What is the volume of the cube?

OpenStudy (anonymous):

sorry i mean it to say 15 centameters

OpenStudy (anonymous):

V=s^3 where s is one edge v=(15)^3